OpenStudy (anonymous):

HELP ME PLEASEEE?!?!?!?!?? FAN & MEDAL!! Polygon ABCD is dilated, rotated, and translated to form polygon A′B′C′D′. The endpoints of AB are at (0, -7) and (8, 8), and the endpoints of A'B' are at (6, -6) and (2, 1.5). What is the scale factor of the dilation?

OpenStudy (anonymous):

@CO_oLBoY

OpenStudy (anonymous):

sorry i don't know how to solve this one

OpenStudy (anonymous):

@ganeshie8 @thomaster @amistre64 @ParthKohli

OpenStudy (anonymous):

xp

OpenStudy (amistre64):

assuming a dilation at the origin the dilation factor d, dilates a point (x,y) to the point (dx,dy)

OpenStudy (amistre64):

but as i see it, we are given the end point after everything else is done ... so we need to rely on length here

OpenStudy (amistre64):

what is the distance of each line segment: AB and A'B'?

OpenStudy (anonymous):

The scale factor of this dilation is 12. It is the ratio of the length of A'B' to the length of AB. soo... yeahh.

OpenStudy (amistre64):

yep length ratios :)

OpenStudy (amistre64):

(0, -7) and (8, 8) ::: (6, -6) and (2, 1.5) +7 +7 -6 6 -6 6 -------------- ---------------- 8,15 4,7.5 sqrt(8^2+15^2) = dsqrt(4^2+7.5^2) sqrt(8^2+15^2) -------------- = d sqrt(4^2+7.5^2)

OpenStudy (anonymous):

Ohh.. hmm..

OpenStudy (amistre64):

hmm, and i did a slight error its spose to be dA = A' not A = dA' so my fraction is upside down

OpenStudy (amistre64):

A' to A is 2 ... but we want A to A' ... so flip it to get 1/2

OpenStudy (amistre64):

yeah AB = 17 A'B' = 8.5 = 17/2

OpenStudy (anonymous):

okay! Hahaa so is that the answer?

OpenStudy (amistre64):

you should be able to deduce it from the postings ... tell me what you have determined the solution to be